OpenStudy (anonymous):

Please help, pleasee! x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (solomonzelman):

brrr... hold on ! what's the question and what are the options ?

OpenStudy (anonymous):

Oh crap sorry I didn't realize the question wasn't on there haha its What is the equation of the line that passes through the points (-2, 3) and (2, 7)?

OpenStudy (solomonzelman):

Can you find the slope ? \(\Huge\color{blue}{ \bf slope~=~\frac{y_{1}-y_{2}}{x_{1}-x_{2}} }\) plug in your points.

OpenStudy (anonymous):

3-7 over -2-2?

OpenStudy (solomonzelman):

\[\frac{3-7 }{-2-2}=\frac{-4 }{-4}=1\]

OpenStudy (anonymous):

Okay what happens after?

OpenStudy (solomonzelman):

Now, you can write y=x+b b is the unknown y-intercept. I am going to plug in one of your given pints to solve for this intercept. y=x+b plunging in (2, 7) (7)=(2)+b 7=2+b b=5 So y = x+5 if you rearrange it , you get -x -x y -x = 5 divide all this by -1, -y+x=5 which is equal to x - y = - 5

OpenStudy (anonymous):

So the answer is C?

OpenStudy (anonymous):

Okay that makes sense now, thank you so much!!!!

OpenStudy (solomonzelman):

You welcome !